OpenStudy (anonymous):

Your school drama club is putting on a play and hopes to raise $600 from ticket sales. They sell adult tickets for $8 each and student tickets for $5 each. Let x represent the number of adult tickets and y represent the number of student tickets. The equation to represent the number of tickets sold to reach their goal, 8x+ 5y = 600, is graphed below. If 40 adult tickets are sold, how many student tickets must be sold to reach their goal?

OpenStudy (ganpat):

adult ticket = x student ticket = y 8x + 5y = 600 given , x = 40 8 (40) + 5y = 600 5y = 600 - 320 5y = 280 y = 56... Students tickets sold...

OpenStudy (jackellyn):

8x+5y=600 x=40 so substitute 40 in the first eqn and solve for y 8(4)+5y=600 ....